# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v1

@@ Abstract
Large Language Models (LLMs) rely on safety alignment to produce socially acceptable responses. This is typically achieved through instruction tuning and reinforcement learning from human feedback. However, this alignment is known to be brittle: further fine-tuning, even on benign or lightly contaminated data, can degrade safety and reintroduce harmful behaviors. A growing body of work suggests that alignment may correspond to identifiable geometric directions in weight space, forming subspaces that could, in principle, be isolated or preserved to defend against misalignment. In this work, we conduct a comprehensive empirical study of this geometric perspective. We examine whether safety-relevant behavior is concentrated in specific subspaces, whether it can be separated from general-purpose learning, and whether harmfulness arises from distinguishable patterns in internal representations. Across both parameter and activation space, our findings are consistent: subspaces that amplify safe behaviors also amplify unsafe ones, and prompts with different safety implications activate overlapping representations. We find no evidence of a subspace that selectively governs safety. These results challenge the assumption that alignment is geometrically localized. Rather than residing in distinct directions, safety appears to emerge from entangled, high-impact components of the model’s broader learning dynamics. This suggests that subspace-based defenses may face fundamental limitations and underscores the need for alternative strategies to preserve alignment under continued training. We corroborate these findings through multiple experiments on five open-source LLMs. Our code is publicly available at: https://github.com/CERT-Lab/safety-subspaces.
@@ 1 Introduction
Large Language Models (LLMs) have shown strong performance across a wide range of general-purpose tasks, including complex reasoning and problem solving [43, 1, 51, 49, 58, 50]. To ensure these models behave responsibly and align with human values, they undergo an additional process of security alignment. Despite known jailbreak methods that can bypass safeguards, aligned models are generally considered significantly safer than their base versions [37, 55, 44]. This alignment is typically achieved through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF), enabling models only improve response quality, and also refuse harmful or inappropriate prompts. A growing line of research focuses on the weight difference between the base and aligned models, commonly referred to as the alignment matrix, which captures the transition from unaligned to aligned behavior. This difference has been used to interpret alignment mechanisms and develop defenses against adversarial attacks [3, 18, 57, 10, 30, 25, 28].
However, this alignment is fragile. Since safety is encoded in the model’s weights, any modification, such as further fine-tuning (FT), can compromise it. While FT adapts models to new tasks by learning update directions, it offers no guarantee that safety is preserved. This exposes a deeper attack surface beyond prompt engineering: an adversary could insert a small number of malicious samples into a training set to subvert alignment [59, 60, 4, 64]. Recent work shows that even benign FT, low-rank adaptation, or pruning can degrade a model’s safety profile [29, 42, 15, 16, 54]. Preserving alignment under continued training is therefore both a practical concern and a theoretically challenging problem. Given this vulnerability, a natural question arises: Does there exist any subspace, whether in weight space or activation space, that uniquely encodes safety alignment? If safety is a distinct and structured property of the model, then updates or representations affecting it might consistently concentrate in identifiable geometric regions. This motivates a broader question: can we isolate or characterize safety-relevant subspaces that amplify aligned behavior or suppress harmful outputs?
To explore this question, we conduct four experiments probing the geometry of safety-related behavior across model weights and internal representations. We begin by analyzing FT updates from purely useful and harmful datasets, projecting them into subspaces derived from the alignment matrix to test whether safety correlates with energy or behavioral impact. Next, we examine contaminated FT, where a small fraction of harmful samples is mixed into a benign dataset. By projecting updates into the orthogonal complement of alignment-derived subspaces, we test whether harmful components can be selectively removed (see Figure 1). In the third experiment, we directly compare the dominant subspaces of useful, harmful, and alignment updates to assess whether safety-altering updates share consistent structure. Finally, we inspect the representation space, comparing internal activations from useful and harmful prompts to ask whether safety-related inputs occupy distinct subspaces, even when weight updates do not.
@@
Across all experiments, we observe a consistent and surprising result: no subspace, whether defined by alignment directions, update energies, or input representations, captures safety-specific behavior in isolation. While certain subspaces, such as the top alignment directions, are behaviorally impactful, they amplify both helpful and harmful behaviors equally, reflecting general update sensitivity rather than alignment. Similarly, activations from harmful and helpful prompts occupy overlapping regions of representation space, offering no evidence for distinct "safety activation" geometry. These findings point to a fundamental limitation of subspace-based alignment strategies. If safe and unsafe behaviors cannot be cleanly separated geometrically, then projection- or filtering-based defenses are unlikely to suppress harmfulness without incurring equivalent losses in utility. Our key contributions are:
@@
  • We show that subspaces derived from alignment updates are not safety-specific; they amplify both helpful and harmful behaviors equally, reflecting general update sensitivity rather than alignment.
@@
  • We find that orthogonal projection intended to filter harmful updates leads to proportional losses in utility, suggesting no selective geometric separation between safe and unsafe behavior.
  • We demonstrate that harmful and aligned updates do not share a consistent subspace, and that harmful prompts do not activate distinct regions of representation space.
@@
  • Through consistent results across five open-source LLMs evaluated in multiple experiments, we challenge the view that safety alignment is geometrically localized and reveal fundamental limitations of subspace-based defenses.
@@ 2 Preliminaries
Let $\mathbf{W}_{0}$ denote the parameters of the base model, and let $\mathbf{W}_{\mathrm{A}}$ represent the parameters of the aligned instruction-tuned model. We further fine-tune the aligned model on a task-specific dataset $\mathcal{D}j$, where $j\in\{\text{Useful},\text{Harmful},\text{Contaminated}\}$, resulting in parameters $\mathbf{W}_{\mathrm{FT},j}$. We decompose the total parameter update as the sum of two components:
[…]
Importance of Alignment Directions ($\Delta_{\mathrm{A}}$).
@@
Alignment training typically emphasizes behavioral properties such as harmlessness, helpfulness, and honesty. Empirical studies [37, 18, 10] suggest that the alignment update $\Delta_{\mathrm{A}}$ encodes directions in parameter space that are strongly correlated with these safety attributes. Our goal is to systematically control the extent to which the subsequent task-specific update $\Delta_{\mathrm{T}}^{j}$ interacts with these alignment directions.
@@
Figure 1: The base model $W_{o}$ is instruction-tuned to produce the aligned model $W_{A}$. Step 1: The difference $\Delta_{A}=W_{A}-W_{o}$ defines a safety direction, from which projection matrices $P_{k}$ (top-K subspace) and $P_{k}^{\perp}$ (orthogonal subspace) are derived. $W_{A}$ is then fine-tuned on three datasets: helpful, harmful, and contaminated, to yield $W_{\text{useful}}$, $W_{\text{harmful}}$, and $W_{\text{contaminated}}$, with updates $\Delta_{t_{j}}$. Step 2: Project $\Delta_{t_{j}}$ using $P_{k}$ and $P_{k}^{\perp}$, and add back to $W_{A}$ to obtain projected models for evaluation. In addition, SVD is performed on the task-specific updates and MSO is computed between the top-K singular vectors.
Constructing the Alignment Subspace.
@@
To formalize this notion, we begin by constructing the alignment subspace. Each tensor in the alignment update $\Delta_{\mathrm{A}}$ is first reshaped into a matrix (flattened if needed) $V_{\mathrm{A}}\in\mathbb{R}^{M\times N}$. We perform a thin singular value decomposition (SVD) of the form $V_{\mathrm{A}}=U\Sigma V^{\top}$, which reveals the principal directions of parameter change [12, 36], ranked by their contribution to the Frobenius norm. The top $k$ (Top-K) right singular vectors in $V$ are then selected to define the alignment subspace:
@@
$\mathcal{S}_{k}:=\operatorname{span}(U_{k}),\quad U_{k}\in\mathbb{R}^{M\times k},\quad k\leq\operatorname{rank}(V_{\mathrm{A}}).$ (3)
@@
Intuitively, $\mathcal{S}_{k}$ captures the $k$ most significant directions of parameter shifts resulting from alignment training. The alignment subspace naturally induces projection operators:
@@
where $P_{k}$ projects a matrix onto the alignment subspace, and $P_{k}^{\perp}$ onto its orthogonal complement.
@@
[…]
Equation 5 retains the update components that align with the alignment directions, while Equation 6 removes this aligned component, retaining only the update orthogonal to the alignment subspace.
@@
To further assess the specificity and effectiveness of the alignment subspace, we introduce two control experiments:
  • Random-K: Instead of using the top-$k$ singular vectors from the SVD of $V_{\mathrm{A}}$, we randomly sample $k$ singular vectors from the full set to construct a randomized alignment subspace.
  • Random: We replace $V_{\mathrm{A}}$ with a random matrix of the same dimensions, perform its SVD, and use the top-$k$ singular vectors to define a synthetic alignment subspace.
@@
We introduce the fractional energy metric to quantify the extent of overlap between the task update and the alignment subspace:
@@
Let $\mathbf{V}\in\mathbb{R}^{d\times n_{V}}$ and $\mathbf{W}\in\mathbb{R}^{d\times n_{W}}$ be two matrices with a shared ambient dimension $d$ but possibly different column counts. We extract their principal directions using thin SVD:
@@
To quantify the similarity between these $\eta$-energy subspaces, we use MSO metric:
@@
Throughout our work, we evaluate both base and instruction-aligned versions of several open-source LLMs. For example, we consider Qwen-2.5 3B (base) alongside its aligned variant, Qwen-2.5 3B Instruct. We report results for base and aligned versions of five models: LLaMA 3.2 1B [11], LLaMA 2 7B [51], Qwen-2.5 1B [58], Qwen-2.5 3B, and Qwen-2.5 7B.
@@ 3 Do Alignment Subspaces Encode Safety?
A central question in understanding safety alignment is whether specific directions in weight space, such as those defined by the difference between a base model and its RLHF-aligned counterpart, encode information unique to safety. If this is the case, then constraining FT updates to lie within these subspaces could offer a principled way to guard against harmful optimization. We begin our investigation by examining whether task-specific FT updates align differently with the top directions of the alignment matrix, depending on whether the task is helpful or harmful.
@@
We fine-tune an aligned instruction-tuned model on two distinct datasets. The first is a 20K subset of MetaMathQA [63], a benchmark of math word problems representing a useful task without safety concerns. The second is a 4K unsafe subset of BeaverTails [26], a synthetic dataset of harmful instruction–response pairs designed to elicit unsafe behavior. We denote the resulting weight updates as $\Delta_{\mathrm{T}}^{\text{Useful}}$ and $\Delta_{\mathrm{T}}^{\text{Harmful}}$ respectively. To quantify behavioral effects, we evaluate harmfulness using the AdvBench dataset [68], with GPT-4o-mini [24] scoring each response from 1 (least harmful) to 5 (most harmful); the final score is the average across samples. Utility is measured by accuracy on the GSM8k test set [8], using final answer correctness. We compute these metrics, energy-kept ratio, utility, and harmfulness, for the projected models $\mathbf{W}_{\text{parallel}}$ and $\mathbf{W}_{\text{orthogonal}}$, as well as for the base, aligned, fine-tuned, and control models.
@@
As shown in Figure 2, the fraction of energy retained in projected updates increases linearly with subspace rank and is consistent across all three subspace types. This pattern holds for both helpful and harmful updates. There is no evidence that update energy is preferentially concentrated in the top directions of $\Delta_{A}$ for safe vs unsafe FT. This suggests that if a "safety subspace" exists, it is not captured simply by energetic alignment with the dominant directions of $\Delta_{A}$. However, while energy is evenly distributed, behavioral impact is not. We can observe that projecting $\Delta_{\mathrm{T}}^{\text{Useful}}$ onto the top-$k$ directions consistently improves utility relative to random projections with equal energy, in Figure 2 and Table 1. Similarly, projecting $\Delta_{\mathrm{T}}^{\text{Harmful}}$ onto the same directions increases harmfulness. Thus, the top singular directions of $\Delta_{A}$ are not uniquely aligned with safety, but they are generally potent. Updates along these directions are more effective, whether the goal is to enhance utility or to elicit harmful behavior. We present results on all models in Table 3 (Appendix B).
@@
Implications: Alignment Directions Reflect Update Sensitivity, Not Safety.
This symmetry across tasks is important. The fact that top-$k$ directions amplify both helpful and harmful behavior equally suggests they do not encode alignment directly. Instead, they appear to represent axes of general parameter sensitivity, directions where updates tend to induce large changes in model behavior. In this sense, $\Delta_{A}$ captures a general learning geometry: directions that are especially effective for optimization, not inherently safe. We draw three key takeaways. First, neither helpful nor harmful updates preferentially align with the top subspaces of $\Delta_{A}$ in terms of energy. Second, those same subspaces are more behaviorally expressive, enhancing both utility and harmfulness depending on the task. Third, this challenges the notion that $\Delta_{A}$ encodes safety-specific information. Its dominant directions support effective learning broadly, without guiding its ethical character. Thus, using $\Delta_{A}$ to constrain updates may regulate the magnitude of behavior change, but not its direction or valence.

### v2

# Safety Subspaces are Not Linearly Distinct: A Fine-Tuning Case Study

@@ Abstract
Large Language Models (LLMs) rely on safety alignment to produce socially acceptable responses. However, this behavior is known to be brittle: further fine-tuning, even on benign or lightly contaminated data, can degrade safety and reintroduce harmful behaviors. A growing body of work suggests that alignment may correspond to identifiable directions in weight space, forming subspaces that could, in principle, be isolated or preserved to defend against misalignment. In this work, we conduct a comprehensive empirical study of this perspective. We examine whether safety-relevant behavior is concentrated in specific linear subspaces, whether it can be separated from general-purpose learning, and whether harmfulness arises from distinguishable patterns in activations. Across both weight and activation spaces, our findings are consistent: subspaces that amplify safe behaviors also amplify useful ones, and prompts with different safety implications activate overlapping representations. Rather than residing in distinct directions, we show that safety is highly entangled with the general learning components of the model. This suggests that subspace-based defenses face fundamental limitations and underscores the need for alternative strategies to preserve safety under continued training. We corroborate these findings with multiple experiments on five open-source LLMs from the Llama and Qwen families. Our code is publicly available at: https://github.com/CERT-Lab/safety-subspaces.
@@ 1 Introduction
Large Language Models (LLMs) show strong performance across a wide range of general-purpose tasks (47, 1, 55, 53, 62, 54). To ensure these models behave responsibly and align with human values, they undergo an additional process of safety alignment. This alignment is typically achieved during the post-training stage, enabling models to improve response quality, and refuse harmful prompts over the pre-trained stage. Despite known jailbreak methods that can bypass safeguards, aligned models are generally considered significantly safer than their base versions (40, 59, 48).
However, this alignment is fragile. Since safety is encoded in the model’s weights, any modification, such as further fine-tuning (FT), can compromise it. This exposes a deeper attack surface beyond prompt engineering: an adversary could insert a small number of malicious samples into a training set to subvert alignment (63, 64, 4, 68). Recent work shows that even benign FT, low-rank adaptation, or pruning can degrade a model’s safety profile (32, 46, 16, 17, 58). A growing line of research seeks to leverage the learned safety adherence to design defenses against adversarial attacks and interpret alignment mechanisms(3, 19, 61, 11, 33, 26, 30).
Fine-tuning (FT) is used to adapt an LLM to new and personalized domains and plays a key role in the widespread adoption of LLMs across diverse contexts. Preserving alignment in this setting, while retaining the improvements achieved through FT, is therefore both a practical concern and a technically challenging problem. This raises a natural question: Does there exist a subspace, whether in weight space or activation space, that uniquely encodes safety alignment information without affecting performance? If such a property exists, it could, in principle, enable the preservation of safety while maintaining model performance under continued training.
To construct defenses, prior works (33, 19) have typically derived safety subspaces using one of two approaches: weight updates from general alignment (aligned–base model deltas) or updates from targeted safety tuning (safety–base model deltas). Our goal is to comprehensively investigate these so-called “safety subspaces” in order to determine whether the information they contain is truly specific to safety. If so, we could separate unsafe information from the useful knowledge learned during FT through simple projections, thereby ensuring that our fine-tuned models are both robust to safety and high-performing.
To explore this question, we design four experiments probing the geometry of safety-related behavior across both model weights and activations. We begin by analyzing FT updates derived from purely useful and purely harmful datasets. These updates are projected into the candidate “safety subspaces” to test whether harmful updates are more expressive than useful ones within these subspaces. Next, we design an experiment involving contaminated FT, where a small fraction of harmful samples is mixed into an otherwise benign dataset. By projecting updates into the orthogonal complement of the candidate subspaces, we test whether harmful components can be selectively removed. From these experiments, we conclude that the candidate subspaces are not safety-specific but instead capture general learning. This leads us to ask whether any distinct safety subspace exists at all. To address this, our third experiment performs pairwise comparisons among useful, harmful, and safety updates to determine which pairs share the greatest similarity. Surprisingly, the similarity between harmful and safety updates is never the highest, as one might expect, and is sometimes even the lowest. Finally, in our fourth experiment, we extend this analysis to activation space, examining whether safety-specific attributes are distinguishable in activations rather than weights.
@@
Across all experiments, we uncover a consistent and surprising result: no subspace, whether in weight or activation space, captures safety-specific behavior in isolation. Although certain subspaces, such as those derived from alignment or safety-specific updates, are impactful, they amplify both safe and useful behaviors alike, indicating that safety is deeply entangled with general learning. Similarly, activations from harmful and helpful prompts occupy overlapping regions of activation space, providing no evidence for distinct safety-related regions. Together, these findings reveal a fundamental limitation of linear subspace-based strategies. Since safe and harmful behaviors cannot be cleanly separated linearly, then projection- or filtering-based defenses are unlikely to suppress harmfulness without incurring comparable losses in utility. Our key contributions are as follows:
@@
  • We show that subspaces derived from alignment and safety-specific updates are not uniquely tied to safety; instead, they amplify both useful and harmful behaviors alike, implying that safety is deeply entangled with general learning (Section 3).
@@
  • We find that orthogonal projection intended to filter harmful updates leads to proportional losses in utility, suggesting no selective linear separation between safe and useful behavior (Section 4).
  • We demonstrate that safety and harmful updates share no relatively significant subspace overlap, confirming that no region of weight space can be isolated specifically for safety (Section 5).
  • Finally, we reaffirm this hypothesis in activation space, showing that harmful prompts do not activate distinct linear regions (Section 6).
@@
  • Through multiple experiments across five open-source LLMs from the Llama and Qwen families, we establish that safety alignment is not linearly separable, revealing fundamental limitations of subspace-based defenses.
@@ 2 Preliminaries
Let $\mathbf{W}_{0}$ denote the parameters of the base model and $\mathbf{W}_{\mathrm{A}}$ that of the aligned model. We denote the parameters of the model after safety tuning, i.e., fine-tuning specifically for harmless responses and refusals, using $\mathbf{W}_{\mathrm{S}}$. We further fine-tune the aligned and the safety-tuned models on a task-specific dataset $\mathcal{D}j$, where $j\in\{\text{Useful},\text{Harmful},\text{Contaminated}\}$, resulting in parameters $\mathbf{W}_{\mathrm{FT},j}$. We decompose the total parameter update as the sum of two components:
$\Delta_{\mathrm{A}}$ $:=\mathbf{W}_{\mathrm{A}}-\mathbf{W}_{0}$ $\Delta_{\mathrm{S}}:=\mathbf{W}_{\mathrm{S}}-\mathbf{W}_{0}$ (alignment/safety-specific updates) (1)
$\Delta_{\mathrm{T}}^{j}$ $:=\mathbf{W}_{\mathrm{FT},j}-\mathbf{W}_{\mathrm{A}}/\mathbf{W}_{\mathrm{S}}$ $\text{(task-specific updates)}.$ (2)
Importance of Alignment Update ($\Delta_{\mathrm{A}}$).
@@
Alignment training typically emphasizes behavioral properties such as harmlessness, helpfulness, and honesty. Empirical studies (40, 19, 11) suggest that the alignment update $\Delta_{\mathrm{A}}$ encodes directions in parameter space that are strongly correlated with these safety attributes. This stage is also the sole point in production model training where safety is explicitly introduced into the model. Our goal is to systematically control the extent to which the subsequent task-specific update $\Delta_{\mathrm{T}}^{j}$ interacts with these alignment directions.
@@
Figure 1: The base model $W_{0}$ is aligned/safety-tuned to produce the model $W_{A/S}$. Step 1: The difference $\Delta_{A/S}=W_{A/S}-W_{o}$ defines an alignment/safety-specific direction, from which projection matrices $P_{k}$ (top-K subspace) and $P_{k}^{\perp}$ (orthogonal subspace) are derived. $W_{A/S}$ is then fine-tuned on three datasets: helpful, harmful, and contaminated, to yield $W_{\text{useful}}$, $W_{\text{harmful}}$, and $W_{\text{contaminated}}$, with updates $\Delta_{t_{j}}$. Step 2: Project $\Delta_{t_{j}}$ using $P_{k}$ and $P_{k}^{\perp}$, and add back to $W_{A/S}$ to obtain projected models for evaluation. In addition, SVD is performed on the task-specific updates, and the Mode Subspace Overlap (MSO) is computed between the top-K singular vectors.
Importance of Safety-Specific Update ($\Delta_{\mathrm{S}}$).
We also aim to capture safety more directly, disentangling it from the broader behavioral changes introduced during the alignment stage. Safety tuning focuses explicitly on refusal and harmlessness, without simultaneously shaping general instruction-following ability. A subtle but important detail is that we use a distribution for safety tuning that is distinct from the one used for harmful fine-tuning, ensuring that our analysis captures genuine interactions rather than artifacts of dataset overlap. Our objective here is to systematically analyze how subsequent task-specific updates $\Delta_{\mathrm{T}}^{j}$ interact with these safety directions.
Constructing the Alignment & Safety Subspaces.
@@
To construct the alignment and safety subspaces, each tensor in the updates $\Delta_{\mathrm{A}}/\Delta_{\mathrm{S}}$ is first reshaped into a matrix (flattened if needed) $V_{\mathrm{A}},V_{\mathrm{S}}\in\mathbb{R}^{M\times N}$. From here on, we use $V_{\mathrm{A/S}}$ to refer to both $V_{\mathrm{A}},V_{\mathrm{S}}$. We perform a thin singular value decomposition (SVD) of the form $V_{\mathrm{A/S}}=U\Sigma V^{\top}$, which reveals the principal directions of parameter change (13, 39), ranked by their contribution to the Frobenius norm. The top $k$ (Top-K) right singular vectors in $V$ are then selected to define the alignment/safety-specific subspace:
@@
$\mathcal{S}_{k}:=\operatorname{span}(U_{k}),\quad U_{k}\in\mathbb{R}^{M\times k},\quad k\leq\operatorname{rank}(V_{\mathrm{A/S}}).$ (3)
@@
Intuitively, $\mathcal{S}_{k}$ captures the $k$ most significant directions of parameter shifts resulting from alignment or safety-specific training. The subspaces naturally induce projection operators:
@@
where $P_{k}$ projects a matrix onto the alignment/safety-specific subspace, and $P_{k}^{\perp}$ onto its orthogonal complement.
@@
$\textbf{Parallel}:$ $\quad\tilde{\Delta}_{\mathrm{T}}^{j}=P_{k}\Delta_{\mathrm{T}}^{j},$ $\quad\mathbf{W}_{\text{parallel}}=\mathbf{W}_{\mathrm{A/S}}+\tilde{\Delta}_{\mathrm{T}}^{j},$ (5)
$\textbf{Orthogonal}:$ $\quad\tilde{\Delta}_{\mathrm{T}}^{j}=P_{k}^{\perp}\Delta_{\mathrm{T}}^{j},$ $\quad\mathbf{W}_{\text{orthogonal}}=\mathbf{W}_{\mathrm{A/S}}+\tilde{\Delta}_{\mathrm{T}}^{j}.$ (6)
Eqn. 5 retains the update components that align with the candidate safety directions, while Eqn. 6 removes this component, retaining only the update orthogonal to the candidate safety subspace. Figure 1 provides an overview of our process.
@@
To further assess the specificity and effectiveness of the chosen safety subspace, we introduce two control experiments:
  • Random-K: Instead of using the top-$k$ singular vectors from the SVD of $V_{\mathrm{A/S}}$, we randomly sample $k$ singular vectors from the full set to construct a randomized safety subspace.
  • Random: We replace $V_{\mathrm{A/S}}$ with a random matrix of the same dimensions, perform its SVD, and use the top-$k$ singular vectors to define a synthetic safety subspace.
@@
We introduce the fractional energy metric to quantify the extent of overlap between the task update and the safety subspace:
@@
Let $\mathbf{V}\in\mathbb{R}^{d\times n_{V}}$ and $\mathbf{W}\in\mathbb{R}^{d\times n_{W}}$ be two matrices with a shared ambient dimension $d$ but possibly different column counts. We extract their principal directions by taking the thin SVD:
@@
To quantify the similarity between these $\eta$-energy subspaces, we use the MSO metric defined as:
@@
We evaluate base and aligned versions of five open-source LLMs: Llama 3.2 1B (12), Llama 2 7B (55), Qwen-2.5 1B (62), Qwen-2.5 3B, and Qwen-2.5 7B. To obtain safety-tuned variants, we fine-tune the base models on the safety-specific BeaverTails dataset (28), using only entries labeled is_safe = True.
@@ 3 Do Alignment Subspaces Encode Safety?
A central question in understanding safety alignment is whether specific directions in weight space, such as those defined by the difference between a base model and its RLHF-aligned counterpart, encode information unique to safety. If this is the case, constraining FT updates to lie within these subspaces could provide a principled approach to guarding against harmful optimization. We begin our investigation by examining whether task-specific FT updates align differently with the top directions of the alignment and safety-specific matrices, depending on whether the task is helpful or harmful.
@@
We fine-tune the aligned and safety-tuned models on two distinct datasets. The first is a 20K subset of MetaMathQA (67), a benchmark of math word problems representing a useful task without safety concerns. The second is a 4K unsafe subset of BeaverTails (28), a synthetic dataset of harmful instruction–response pairs designed to elicit unsafe behavior. The resulting weight updates are denoted as $\Delta_{\mathrm{T}}^{\text{Useful}}$ and $\Delta_{\mathrm{T}}^{\text{Harmful}}$, respectively. To quantify behavioral effects, we evaluate harmfulness on AdvBench (72), with GPT-4o-mini (25) scoring each response from 1 (least harmful) to 5 (most harmful); the final score is the average across samples. Utility is measured as accuracy on the GSM8k test set (9), based on final-answer correctness. For each setting, we compute harmfulness, utility, and the energy-kept ratio for the projected models $\mathbf{W}{{}_{\text{parallel}}}$ and $\mathbf{W}{{}_{\text{orthogonal}}}$, as well as for the base, aligned, fine-tuned, and control models.
@@
As shown in Figures 2 and 6 (Appendix C), the fraction of energy retained in projected updates increases linearly with subspace rank and is consistent across all three subspace types. This pattern holds for both helpful and harmful updates. We find no evidence that update energy is preferentially concentrated in the top directions of $\Delta_{\mathrm{A/S}}$ for safe versus unsafe FT. This suggests that if a “safety subspace” exists, it is not revealed simply through energetic alignment with the dominant directions of $\Delta_{\mathrm{A/S}}$. At the same time, while energy is evenly distributed, behavioral impact is not. Figure 2 and Table 1 show that projecting $\Delta_{\mathrm{T}}^{\text{Useful}}$ onto the top-$k$ directions consistently improves utility relative to random projections with equal energy. Similarly, projecting $\Delta_{\mathrm{T}}^{\text{Harmful}}$ onto the same directions increases harmfulness. Thus, the top singular directions of $\Delta_{\mathrm{A/S}}$ are not uniquely aligned with safety, but they are generally potent: updates along these directions are more effective, whether the goal is to enhance utility or to elicit harmful behavior. Comprehensive results for all models are provided in Table 4 (Appendix C).
@@
Implications: Alignment Directions Reflect General Learning, Not Safety.
This symmetry across tasks is important. The fact that top-$k$ directions amplify both helpful and harmful behaviors equally suggests they do not encode alignment directly. Instead, they represent axes of general parameter sensitivity, ie. directions where updates tend to induce large changes in model behavior. This holds for both alignment and safety-specific updates, implying that disentangling safety offers no clear separation or benefit. In this sense, $\Delta_{\mathrm{A/S}}$ captures a general learning geometry: directions that are highly effective for optimization but not inherently safe. We draw three key takeaways. First, neither helpful nor harmful updates preferentially align with the top subspaces of $\Delta_{\mathrm{A}}$ or $\Delta_{\mathrm{S}}$ in terms of energy. Second, these same subspaces are more behaviorally expressive, amplifying both utility and harmfulness depending on the task. Third, this challenges the assumption that $\Delta_{\mathrm{A/S}}$ encode safety-specific information expressed in their top subspaces. Thus, using $\Delta_{\mathrm{A/S}}$ to constrain updates regulates the magnitude of behavioral change, but not its ethical nature.